3.2c A compound is found to contain Si, Cl, and I with the following percent composition:...

3.2c

A compound is found to contain Si, Cl, and I with the following percent composition:

Si:    7.962%
Cl:    20.10%
I:    71.94%


What is the empirical formula of this compound?


in the order Si, Cl, I.
Answer #1

we have mass of each elements as:

Si: 7.962 g

Cl: 20.1 g

I: 71.94 g

Divide by molar mass to get number of moles of each:

Si: 7.962/28.09 = 0.2834

Cl: 20.1/35.45 = 0.567

I: 71.94/126.9 = 0.5669

Divide by smallest to get simplest whole number ratio:

Si: 0.2834/0.2834 = 1

Cl: 0.567/0.2834 = 2

I: 0.5669/0.2834 = 2

  

So empirical formula is: SiCl2I2

Answer: SiCl2I2